# Hyperbolic nonlinear Schrödinger equations on $\mathbb{R}\times\mathbb{T}$

### v2

@@ 4. Global Strichartz estimates
$\begin{split}\|e^{it\Box}P_{\leq N}f\|_{\ell^{8}L^{4}(\mathbb{R},L^{4}(\mathcal{M}))}&\leq C_{\epsilon}N^{\epsilon}\|f\|_{L^{2}(\mathcal{M})}.\end{split}$ (4.1)
@@
$\begin{split}\|e^{it\Box}P_{\leq N}f\|_{\ell^{4}L^{\infty}(\mathbb{R},L^{\infty}(\mathcal{M}))}&\leq CN\|f\|_{L^{2}(\mathcal{M})}.\end{split}$ (4.2)
@@
$\begin{split}\|e^{it\Box}P_{\leq N}f\|_{\ell^{q}L^{p}(\mathbb{R},L^{p}(\mathcal{M}))}&\leq C_{\epsilon}N^{1-\frac{4}{p}+\epsilon}\|f\|_{L^{2}(\mathcal{M})}.\end{split}$ (4.3)
@@
$\Big\|\int_{\mathbb{R}}e^{-is\Box_{\mathbf{x}}}P_{\leq N}f(s,\mathbf{x})\,ds\Big\|_{L^{2}_{\mathbf{x}}(\mathcal{M})}\lesssim N\|f\|_{\ell^{\frac{4}{3}}_{\gamma}L^{1}_{t,\mathbf{x}}([\gamma,\gamma+1)\times\mathcal{M})}.$ (4.4)
@@
$\begin{aligned} \sigma_{\text{diag}}\leq\sum_{\alpha,|\gamma|\leq 10}\Big\|\int_{\mathbb{R}}e^{-is\Box_{\mathbf{x}}}P_{\leq N}f_{\alpha+\gamma}(s,\mathbf{x})\,ds\Big\|_{L^{2}_{\mathbf{x}}}^{2}\end{aligned}.$ (4.6)
@@
$\Big\|\int_{\mathbb{R}}e^{-it\Box_{\mathbf{x}}}P_{\leq N}h(t,\mathbf{x})\,dt\Big\|_{L^{2}_{\mathbf{x}}(\mathcal{M})}\lesssim N\|h\|_{L^{1}_{t,\mathbf{x}}(I\times\mathcal{M})}.$ (4.7)
@@
$\sigma_{\text{diag}}$ $\lesssim N^{2}\sum_{\alpha,|\gamma|\leq 10}\|f_{\alpha+\gamma}(t)\|_{L^{1}_{t,\mathbf{x}}([-1,1]\times\mathcal{M})}^{2}$ (4.8)
@@
$\sigma_{\text{non-diag}}$ $\lesssim\sum_{\alpha,|\gamma|\geq 10}\|K_{N,\gamma}*_{t,\mathbf{x}}P_{\leq N}f_{\alpha}\|_{L^{\infty}_{t,\mathbf{x}}}\|f_{\alpha+\gamma}\|_{L^{1}_{t,\mathbf{x}}}$ (4.9)
@@
$\mathcal{J}:=\sum_{\gamma\in\mathbb{Z}}J_{\gamma}^{8}\lesssim_{\epsilon}N^{\epsilon},$ (4.10)
@@
$J_{\gamma}:=\|e^{-\frac{1}{4}(t-\gamma)^{2}}e^{it\Box}\mathcal{F}_{\xi,k}\big[\mathds{1}_{\{|\xi|+|k|\lesssim N\}}f\big]\|_{L^{4}_{x,y,t}(\mathbb{R}\times\mathbb{T}\times\mathbb{R})}.$ (4.11)
@@
$\begin{split}{\vec{\xi}}&=(\xi_{1},\xi_{2},\xi_{3},\xi_{4}),\qquad\quad{\vec{k}}=(k_{1},k_{2},k_{3},k_{4}),\qquad\quad\vec{\iota}=(1,-1,1,-1),\\ \vec{\xi}\cdot\vec{\iota}&=\xi_{1}-\xi_{2}+\xi_{3}-\xi_{4},\qquad\vec{k}\cdot\vec{\iota}=k_{1}-k_{2}+k_{3}-k_{4},\\ f_{j}&=\mathds{1}_{\{|\xi_{j}|+|k_{j}|\lesssim N\}}f(\xi_{j},k_{j}),\,\,j\in\{1,3\},\qquad f_{j}=\mathds{1}_{\{|\xi_{j}|+|k_{j}|\lesssim N\}}\overline{f}(\xi_{j},k_{j}),\,\,j\in\{2,4\},\\ Q(\xi,k)&=|(\xi_{1},k_{1})|^{2}_{-}+|(\xi_{3},k_{3})|^{2}_{-}-|(\xi_{2},k_{2})|^{2}_{-}-|(\xi_{4},k_{4})|^{2}_{-}\end{split}$ (4.12)
@@
$\begin{split}\mathcal{J}&=16\pi^{5}\sum_{\begin{subarray}{c}k_{1}\dots k_{4}\\ k^{\prime}_{1}\dots k^{\prime}_{4}\end{subarray}}\int_{\begin{subarray}{c}\xi_{1}\dots\xi_{4}\\ \xi_{1}^{\prime}\dots\xi_{4}^{\prime}\end{subarray}}\Pi_{j=1}^{4}f_{j}\overline{\Pi_{l=1}^{4}f_{l}^{\prime}}\,e^{-\frac{1}{4}\left[Q^{2}+(Q^{\prime})^{2}\right]}\Big(\sum_{\gamma}e^{-i\gamma\left[Q-Q^{\prime}\right]}\Big)\\ &\qquad\qquad\quad\times\delta(\vec{\xi}\cdot\vec{\iota})\delta(\vec{\xi^{\prime}}\cdot\vec{\iota})\delta(\vec{k}\cdot\vec{\iota})\delta(\vec{k^{\prime}}\cdot\vec{\iota})d{\vec{\xi}}d{\vec{\xi}^{\prime}}\\ &=16\sqrt{2}\pi^{\frac{11}{2}}\sum_{\begin{subarray}{c}k_{1}\dots k_{4}\\ k^{\prime}_{1}\dots k^{\prime}_{4}\end{subarray}}\int_{\begin{subarray}{c}\xi_{1}\dots\xi_{4}\\ \xi_{1}^{\prime}\dots\xi_{4}^{\prime}\end{subarray}}\Pi_{j=1}^{4}f_{j}\overline{\Pi_{l=1}^{4}f_{l}^{\prime}}\,e^{-\frac{1}{4}\left[Q^{2}+(Q^{\prime})^{2}\right]}\\ &\qquad\quad\quad\times\Big(\sum_{\mu\in 2\pi\mathbb{Z}}\delta(\mu-Q+Q^{\prime})\Big)\delta(\vec{\xi}\cdot\vec{\iota})\delta(\vec{\xi^{\prime}}\cdot\vec{\iota})\delta(\vec{k}\cdot\vec{\iota})\delta(\vec{k^{\prime}}\cdot\vec{\iota})d{\vec{\xi}}d{\vec{\xi}^{\prime}}\end{split}$ (4.13)
@@
$\begin{split}\sum_{|k|,|k^{\prime}|\lesssim N}\int_{|\xi|,|\xi^{\prime}|\lesssim N}&e^{-\frac{1}{2}\left[||(\xi,k)-{\vec{C}}|^{2}_{-}-|R|_{-}^{2}|^{2}+||(\xi^{\prime},k^{\prime})-{\vec{C}}^{\prime}|^{2}_{-}-|R^{\prime}|_{-}^{2}|^{2}\right]}\\ &\times\delta(|(\xi,k)-{\vec{C}}|^{2}_{-}+|(\xi^{\prime},k^{\prime})-{\vec{C}^{\prime}}|^{2}_{-}-A)d\xi d\xi^{\prime}\lesssim(\log N)^{2}\end{split}$ (4.15)
@@
$A_{j}=\{(\xi,k)\in\mathbb{R}\times\mathbb{Z}:|\Phi(\xi,k)|\in[j,j+1),\,|\xi|+|k|\lesssim N\},\quad j\geq 0,$ (4.16)
@@
$\sup_{\begin{subarray}{c}|a|,|b|\lesssim N\end{subarray}}{\rm mes}(\{(\xi,k)\in\mathbb{R}\times\mathbb{Z}:C_{0}\leq(\xi-a)^{2}-(k-b)^{2}\leq C_{0}+K,\,|\xi|+|k|\lesssim N\})\lesssim K\log N$ (4.19)
@@
$\begin{split}\sup_{A\in\mathbb{R}}\iint_{\mathbb{R}^{2}}\delta(\zeta^{2}+\eta^{2}-A)d\zeta d\eta&=\pi.\end{split}$ (4.20)
@@
${\bf I}_{jp}^{1}\lesssim\sup_{B\in\mathbb{R}}\iint_{\mathbb{R}^{2}}\delta(\xi^{2}+(\xi^{\prime})^{2}-B)d\xi d\xi^{\prime}\sum_{k,k^{\prime}=O(1)}1\lesssim 1.$ (4.21)
@@
${\bf I}_{jp}^{2}$ $\lesssim\Big(\sup_{\begin{subarray}{c}B\in\mathbb{R}\\ C\in\mathbb{R},c\in\frac{1}{2}\mathbb{Z}\end{subarray}}\sum_{|k|\lesssim N}\int_{\mathbb{R}}\mathds{1}_{\big[\sqrt{|k-c|^{2}+C},\sqrt{|k-c|^{2}+C+1}\big)}(\xi)\delta(\xi^{2}-B)d\xi\Big)\sum_{k^{\prime}}\int_{\mathbb{R}}\mathds{1}_{A_{p}^{\prime}}(\xi^{\prime},k^{\prime})d\xi^{\prime}$ (4.22)
@@
$\|\mathds{1}_{I}\cdot P_{C}u\|_{\ell_{\gamma}^{q}L^{p}([\gamma,\gamma+1]\times\mathbb{R}\times\mathbb{T})}\lesssim N^{1-\frac{4}{p}}\|P_{C}u\|_{U^{\min\{p,q\}}_{\Box}(I;L^{2})}\lesssim N^{1-\frac{4}{p}}\|P_{C}u\|_{Y^{0}(I)},$ (5.1)
@@
$\Big|\int_{I}\int_{\mathcal{M}}\prod_{l=1}^{2k+1}\widetilde{u}_{l}\overline{v}\,dxdt\Big|\lesssim\|v\|_{Y^{-s}(I)}\|u_{1}\|_{Y^{s}(I)}\prod_{l=2}^{2k+1}\|u_{l}\|_{Y^{s_{2,k}}(I)}$ (5.2)
@@
$4<p<\frac{8k}{k+1}\quad\text{for}\,\,k\geq 3,$ (5.3)
@@
$\|\mathds{1}_{I}\cdot P_{N}u\|_{\ell^{q_{2,k}}L^{p_{2,k}}([\gamma,\gamma+1]\times\mathcal{M})}\lesssim N^{s_{2,k}}\|P_{N}u\|_{Y^{0}(I)}.$ (5.4)
@@
$\frac{2}{p}+\frac{k-2}{p_{2,k}}+\frac{1}{\widetilde{p}}=\frac{1}{2},$ (5.5)
@@
$1-\frac{4}{\widetilde{p}}-s_{2,k}=-2+\frac{8}{p}+s_{2,k}=\delta>0.$ (5.6)
@@
$\Big|\int_{I}\int_{\mathcal{M}}\prod_{l=1}^{2k+1}\widetilde{u}_{l}\overline{v}\,dxdydt\Big|\lesssim\sum_{N_{0}\sim N_{1}\geq\cdots\geq N_{2k+1}}\Big\|P_{N_{0}}v\prod_{l=1}^{2k+1}P_{N_{l}}u_{l}\Big\|_{L^{1}_{t,x,y}}$ (5.7)
@@
$\Big\|P_{C_{\alpha}}P_{N_{0}}v\prod_{l=1}^{k}P_{N_{2l}}u_{2l}\Big\|_{L^{2}_{t,x,y}}^{2}=\sum_{\gamma\in\mathbb{Z}}\Big\|P_{C_{\alpha}}P_{N_{0}}v\prod_{l=1}^{k}P_{N_{2l}}u_{2l}\Big\|_{L^{2}_{t,\mathbf{x}}(I_{\gamma}\times\mathcal{M})}^{2}$ (5.8)
@@
$\Big\|P_{C_{\beta}}P_{N_{1}}u_{1}\prod_{l=1}^{k}P_{N_{2l+1}}u_{2l+1}\Big\|_{L^{2}_{t,x,y}}\lesssim N_{3}^{2-\frac{8}{p}-s_{2,k}}N_{2k+1}^{1-\frac{4}{\widetilde{p}}-s_{2,k}}\|P_{C_{\beta}}P_{N_{1}}u_{1}\|_{Y^{0}}\prod_{l=1}^{k}\|P_{N_{2l+1}}u_{2l+1}\|_{Y^{s_{2,k}}}.$ (5.9)
@@
$\Big\|\int_{0}^{t}e^{i(t-\tau)\Box}\prod_{j=1}^{2k+1}\widetilde{u}_{j}\,d\tau\Big\|_{X^{s}(I)}\lesssim\sum_{l=1}^{2k+1}\|u_{l}\|_{X^{s}(I)}\prod_{\begin{subarray}{c}j=1\\ j\neq l\end{subarray}}^{2k+1}\|u_{j}\|_{X^{s_{2,k}}(I)}$ (5.10)
@@
$\Gamma(u)(t):=e^{it\Box}u_{0}\pm i\mathcal{I}[F(u)](t),\quad t\in I\setminus\{0\},$ (5.11)
@@
$\lim_{t\rightarrow\infty}e^{-it\Box}u(t)=u_{+}\in H^{s_{2,k}}(\mathbb{R}\times\mathbb{T}).$ (5.12)

### v3

@@ 4. Global Strichartz estimates
(4.1) $\begin{split}\|e^{it\Box}P_{\leq N}f\|_{\ell^{8}L^{4}(\mathbb{R},L^{4}(\mathcal{M}))}&\leq C_{\epsilon}N^{\epsilon}\|f\|_{L^{2}(\mathcal{M})}.\end{split}$
@@
(4.2) $\begin{split}\|e^{it\Box}P_{\leq N}f\|_{\ell^{4}L^{\infty}(\mathbb{R},L^{\infty}(\mathcal{M}))}&\leq CN\|f\|_{L^{2}(\mathcal{M})}.\end{split}$
@@
(4.3) $\begin{split}\|e^{it\Box}P_{\leq N}f\|_{\ell^{q}L^{p}(\mathbb{R},L^{p}(\mathcal{M}))}&\leq C_{\epsilon}N^{1-\frac{4}{p}+\epsilon}\|f\|_{L^{2}(\mathcal{M})}.\end{split}$
@@
(4.4) $\Big\|\int_{\mathbb{R}}e^{-is\Box_{\mathbf{x}}}P_{\leq N}f(s,\mathbf{x})\,ds\Big\|_{L^{2}_{\mathbf{x}}(\mathcal{M})}\lesssim N\|f\|_{\ell^{\frac{4}{3}}_{\gamma}L^{1}_{t,\mathbf{x}}([\gamma,\gamma+1)\times\mathcal{M})}.$
@@
(4.6) $\begin{aligned} \sigma_{\text{diag}}\leq\sum_{\alpha,|\gamma|\leq 10}\Big\|\int_{\mathbb{R}}e^{-is\Box_{\mathbf{x}}}P_{\leq N}f_{\alpha+\gamma}(s,\mathbf{x})\,ds\Big\|_{L^{2}_{\mathbf{x}}}^{2}\end{aligned}.$
@@
(4.7) $\Big\|\int_{\mathbb{R}}e^{-it\Box_{\mathbf{x}}}P_{\leq N}h(t,\mathbf{x})\,dt\Big\|_{L^{2}_{\mathbf{x}}(\mathcal{M})}\lesssim N\|h\|_{L^{1}_{t,\mathbf{x}}(I\times\mathcal{M})}.$
@@
(4.8) $\sigma_{\text{diag}}$ $\lesssim N^{2}\sum_{\alpha,|\gamma|\leq 10}\|f_{\alpha+\gamma}(t)\|_{L^{1}_{t,\mathbf{x}}([-1,1]\times\mathcal{M})}^{2}$
@@
(4.9) $\sigma_{\text{non-diag}}$ $\lesssim\sum_{\alpha,|\gamma|\geq 10}\|K_{N,\gamma}*_{t,\mathbf{x}}P_{\leq N}f_{\alpha}\|_{L^{\infty}_{t,\mathbf{x}}}\|f_{\alpha+\gamma}\|_{L^{1}_{t,\mathbf{x}}}$
@@
(4.10) $\mathcal{J}:=\sum_{\gamma\in\mathbb{Z}}J_{\gamma}^{8}\lesssim_{\epsilon}N^{\epsilon},$
@@
(4.11) $J_{\gamma}:=\|e^{-\frac{1}{4}(t-\gamma)^{2}}e^{it\Box}\mathcal{F}_{\xi,k}\big[\mathds{1}_{\{|\xi|+|k|\lesssim N\}}f\big]\|_{L^{4}_{x,y,t}(\mathbb{R}\times\mathbb{T}\times\mathbb{R})}.$
@@
(4.12) $\begin{split}{\vec{\xi}}&=(\xi_{1},\xi_{2},\xi_{3},\xi_{4}),\qquad\quad{\vec{k}}=(k_{1},k_{2},k_{3},k_{4}),\qquad\quad\vec{\iota}=(1,-1,1,-1),\\ \vec{\xi}\cdot\vec{\iota}&=\xi_{1}-\xi_{2}+\xi_{3}-\xi_{4},\qquad\vec{k}\cdot\vec{\iota}=k_{1}-k_{2}+k_{3}-k_{4},\\ f_{j}&=\mathds{1}_{\{|\xi_{j}|+|k_{j}|\lesssim N\}}f(\xi_{j},k_{j}),\,\,j\in\{1,3\},\qquad f_{j}=\mathds{1}_{\{|\xi_{j}|+|k_{j}|\lesssim N\}}\overline{f}(\xi_{j},k_{j}),\,\,j\in\{2,4\},\\ Q(\xi,k)&=|(\xi_{1},k_{1})|^{2}_{-}+|(\xi_{3},k_{3})|^{2}_{-}-|(\xi_{2},k_{2})|^{2}_{-}-|(\xi_{4},k_{4})|^{2}_{-}\end{split}$
@@
(4.13) $\begin{split}\mathcal{J}&=16\pi^{5}\sum_{\begin{subarray}{c}k_{1}\dots k_{4}\\ k^{\prime}_{1}\dots k^{\prime}_{4}\end{subarray}}\int_{\begin{subarray}{c}\xi_{1}\dots\xi_{4}\\ \xi_{1}^{\prime}\dots\xi_{4}^{\prime}\end{subarray}}\Pi_{j=1}^{4}f_{j}\overline{\Pi_{l=1}^{4}f_{l}^{\prime}}\,e^{-\frac{1}{4}\left[Q^{2}+(Q^{\prime})^{2}\right]}\Big(\sum_{\gamma}e^{-i\gamma\left[Q-Q^{\prime}\right]}\Big)\\ &\qquad\qquad\quad\times\delta(\vec{\xi}\cdot\vec{\iota})\delta(\vec{\xi^{\prime}}\cdot\vec{\iota})\delta(\vec{k}\cdot\vec{\iota})\delta(\vec{k^{\prime}}\cdot\vec{\iota})d{\vec{\xi}}d{\vec{\xi}^{\prime}}\\ &=16\sqrt{2}\pi^{\frac{11}{2}}\sum_{\begin{subarray}{c}k_{1}\dots k_{4}\\ k^{\prime}_{1}\dots k^{\prime}_{4}\end{subarray}}\int_{\begin{subarray}{c}\xi_{1}\dots\xi_{4}\\ \xi_{1}^{\prime}\dots\xi_{4}^{\prime}\end{subarray}}\Pi_{j=1}^{4}f_{j}\overline{\Pi_{l=1}^{4}f_{l}^{\prime}}\,e^{-\frac{1}{4}\left[Q^{2}+(Q^{\prime})^{2}\right]}\\ &\qquad\quad\quad\times\Big(\sum_{\mu\in 2\pi\mathbb{Z}}\delta(\mu-Q+Q^{\prime})\Big)\delta(\vec{\xi}\cdot\vec{\iota})\delta(\vec{\xi^{\prime}}\cdot\vec{\iota})\delta(\vec{k}\cdot\vec{\iota})\delta(\vec{k^{\prime}}\cdot\vec{\iota})d{\vec{\xi}}d{\vec{\xi}^{\prime}}\end{split}$
@@
(4.15) $\begin{split}\sum_{|k|,|k^{\prime}|\lesssim N}\int_{|\xi|,|\xi^{\prime}|\lesssim N}&e^{-\frac{1}{2}\left[||(\xi,k)-{\vec{C}}|^{2}_{-}-|R|_{-}^{2}|^{2}+||(\xi^{\prime},k^{\prime})-{\vec{C}}^{\prime}|^{2}_{-}-|R^{\prime}|_{-}^{2}|^{2}\right]}\\ &\times\delta(|(\xi,k)-{\vec{C}}|^{2}_{-}+|(\xi^{\prime},k^{\prime})-{\vec{C}^{\prime}}|^{2}_{-}-A)d\xi d\xi^{\prime}\lesssim(\log N)^{2}\end{split}$
@@
(4.16) $A_{j}=\{(\xi,k)\in\mathbb{R}\times\mathbb{Z}:|\Phi(\xi,k)|\in[j,j+1),\,|\xi|+|k|\lesssim N\},\quad j\geq 0,$
@@
(4.19) $\sup_{\begin{subarray}{c}|a|,|b|\lesssim N\end{subarray}}{\rm mes}(\{(\xi,k)\in\mathbb{R}\times\mathbb{Z}:C_{0}\leq(\xi-a)^{2}-(k-b)^{2}\leq C_{0}+K,\,|\xi|+|k|\lesssim N\})\lesssim K\log N$
@@
(4.20) $\begin{split}\sup_{A\in\mathbb{R}}\iint_{\mathbb{R}^{2}}\delta(\zeta^{2}+\eta^{2}-A)d\zeta d\eta&=\pi.\end{split}$
@@
(4.21) ${\bf I}_{jp}^{1}\lesssim\sup_{B\in\mathbb{R}}\iint_{\mathbb{R}^{2}}\delta(\xi^{2}+(\xi^{\prime})^{2}-B)d\xi d\xi^{\prime}\sum_{k,k^{\prime}=O(1)}1\lesssim 1.$
@@
(4.22) ${\bf I}_{jp}^{2}$ $\lesssim\Big(\sup_{\begin{subarray}{c}B\in\mathbb{R}\\ C\in\mathbb{R},c\in\frac{1}{2}\mathbb{Z}\end{subarray}}\sum_{|k|\lesssim N}\int_{\mathbb{R}}\mathds{1}_{\big[\sqrt{|k-c|^{2}+C},\sqrt{|k-c|^{2}+C+1}\big)}(\xi)\delta(\xi^{2}-B)d\xi\Big)\sum_{k^{\prime}}\int_{\mathbb{R}}\mathds{1}_{A_{p}^{\prime}}(\xi^{\prime},k^{\prime})d\xi^{\prime}$
@@
(5.1) $\|\mathds{1}_{I}\cdot P_{C}u\|_{\ell_{\gamma}^{q}L^{p}([\gamma,\gamma+1]\times\mathbb{R}\times\mathbb{T})}\lesssim N^{1-\frac{4}{p}}\|P_{C}u\|_{U^{\min\{p,q\}}_{\Box}(I;L^{2})}\lesssim N^{1-\frac{4}{p}}\|P_{C}u\|_{Y^{0}(I)},$
@@
(5.2) $\Big|\int_{I}\int_{\mathcal{M}}\prod_{l=1}^{2k+1}\widetilde{u}_{l}\overline{v}\,dxdt\Big|\lesssim\|v\|_{Y^{-s}(I)}\|u_{1}\|_{Y^{s}(I)}\prod_{l=2}^{2k+1}\|u_{l}\|_{Y^{s_{2,k}}(I)}$
@@
(5.3) $4<p<\frac{8k}{k+1}\quad\text{for}\,\,k\geq 3,$
@@
(5.4) $\|\mathds{1}_{I}\cdot P_{N}u\|_{\ell^{q_{2,k}}L^{p_{2,k}}([\gamma,\gamma+1]\times\mathcal{M})}\lesssim N^{s_{2,k}}\|P_{N}u\|_{Y^{0}(I)}.$
@@
(5.5) $\frac{2}{p}+\frac{k-2}{p_{2,k}}+\frac{1}{\widetilde{p}}=\frac{1}{2},$
@@
(5.6) $1-\frac{4}{\widetilde{p}}-s_{2,k}=-2+\frac{8}{p}+s_{2,k}=\delta>0.$
@@
(5.7) $\Big|\int_{I}\int_{\mathcal{M}}\prod_{l=1}^{2k+1}\widetilde{u}_{l}\overline{v}\,dxdydt\Big|\lesssim\sum_{N_{0}\sim N_{1}\geq\cdots\geq N_{2k+1}}\Big\|P_{N_{0}}v\prod_{l=1}^{2k+1}P_{N_{l}}u_{l}\Big\|_{L^{1}_{t,x,y}}$
@@
(5.8) $\Big\|P_{C_{\alpha}}P_{N_{0}}v\prod_{l=1}^{k}P_{N_{2l}}u_{2l}\Big\|_{L^{2}_{t,x,y}}^{2}=\sum_{\gamma\in\mathbb{Z}}\Big\|P_{C_{\alpha}}P_{N_{0}}v\prod_{l=1}^{k}P_{N_{2l}}u_{2l}\Big\|_{L^{2}_{t,\mathbf{x}}(I_{\gamma}\times\mathcal{M})}^{2}$
@@
(5.9) $\Big\|P_{C_{\beta}}P_{N_{1}}u_{1}\prod_{l=1}^{k}P_{N_{2l+1}}u_{2l+1}\Big\|_{L^{2}_{t,x,y}}\lesssim N_{3}^{2-\frac{8}{p}-s_{2,k}}N_{2k+1}^{1-\frac{4}{\widetilde{p}}-s_{2,k}}\|P_{C_{\beta}}P_{N_{1}}u_{1}\|_{Y^{0}}\prod_{l=1}^{k}\|P_{N_{2l+1}}u_{2l+1}\|_{Y^{s_{2,k}}}.$
@@
(5.10) $\Big\|\int_{0}^{t}e^{i(t-\tau)\Box}\prod_{j=1}^{2k+1}\widetilde{u}_{j}\,d\tau\Big\|_{X^{s}(I)}\lesssim\sum_{l=1}^{2k+1}\|u_{l}\|_{X^{s}(I)}\prod_{\begin{subarray}{c}j=1\\ j\neq l\end{subarray}}^{2k+1}\|u_{j}\|_{X^{s_{2,k}}(I)}$
@@
(5.11) $\Gamma(u)(t):=e^{it\Box}u_{0}\pm i\mathcal{I}[F(u)](t),\quad t\in I\setminus\{0\},$
@@
(5.12) $\lim_{t\rightarrow\infty}e^{-it\Box}u(t)=u_{+}\in H^{s_{2,k}}(\mathbb{R}\times\mathbb{T}).$